OpenStudy (anonymous):

Let z=13+7i and w=3[cos(1.43)+isin(1.43)] A) Convert z to polar form B) Calculate zw using De Moivre's theorem C) Calculate z/w using De Moivre's theorem

OpenStudy (anonymous):

For \(z=x+iy=r\exp(i\theta)\), you have \[\begin{cases}r=\sqrt{x^2+y^2}\\\theta=\tan^{-1}\dfrac{y}{x}\end{cases}\] \[r=\sqrt{13^2+7^2}=\sqrt{218}\approx14.765\] \[\theta=\tan^{-2}\frac{7}{13}\approx0.494\]

OpenStudy (anonymous):

Then, for \(z=r\exp(i\theta)\) and \(w=s\exp(i\phi)\), you have \[z\cdot w=rs\exp(i(\theta+\phi))~~~\text{and}~~~\frac{z}{w}=\frac{r}{s}\exp(i(\theta-\phi))\]

OpenStudy (anonymous):

Thanks